# UFID: A Unified Framework for Input-level Backdoor Detection on Diffusion Models

### v1

@@ 3.3 Scenario 2: Conditional Diffusion Models
Conditional diffusion models accept input from other modalities to guide the generation of user-intended images. For instance, in stable diffusion [36], textual input is used to query the model. However, it is evident that the detection approach employed for unconditional diffusion models cannot be directly applied to conditional diffusion models due to the inability to introduce Gaussian noise to discrete textual input. To adapt our detection method for conditional diffusion models, we have extended the detection approach with slight modifications, as depicted in the second row of Figure 2. We leverage variations in output diversity to distinguish between clean and backdoor samples. To further accentuate this diversity gap and enhance distinguishability in conditional setting, we propose appending the input text $x_{i}$ with a random phrase $ph_{j}$ selected from a diverse phrase pool containing completely distinct phrases, such as "Iron Man" and "Kitchen Dish Washer," denoted as =$\mathbb{N}=\{ph_{1},ph_{2},...ph_{|\mathbb{N}|}\}$, where $|\mathbb{M}|\ll|\mathbb{N}|$. For each input $x_{i}$, we repeat this process $|\mathbb{M}|$ times, resulting in augmented an input batch $\mathbb{I}=\{x_{i}\}\cup\{x_{i}^{j}|x_{i}^{j}=x_{i}\oplus ph_{j},\forall 1\leq j\leq|\mathbb{M}|\}$, where $\oplus$ denotes the string appending operator. Similarly, we can generate an image batch $y_{i}$ by querying the target stable diffusion model using Equation 2.
3.4 A Unified Framework for Backdoor Detection
@@
As previously discussed, the diversity of images generated by a specific input $i$ can be leveraged to detect backdoors. To quantify this diversity, we employ a two-step process that involves calculating both pairwise and overall similarities within the generated batch. The detailed steps are as follows:
Pairwise Similarity Calculation.
@@
Initially, we calculate semantic embeddings of generated images through a pre-trained image encoder (e.g., ViT-ImageNet [44] and CLIP [35]), denoted as $f_{\mathcal{E}}(\cdot)$. Subsequently, we calculate the local similarity for each pair of images in the generated batch using cosine similarity, represented by $S_{c}(\cdot,\cdot)$.
Graph Density Calculation.
@@
Following this, we construct a weighted graph and compute its graph density to represent the overall similarity of all images within graph. Specifically, let $G_{i}=(V_{i},\mathcal{E}_{i})$ represent the similarity graph for input sample $x_{i}$, where $|V_{i}|=|\mathbb{M}|$ constitutes the set of vertices (symbolizing the generated images) and $\mathcal{E}_{i}$ is the set of edges. Each edge’s weight, connecting a pair of images $u,v\in V$, indicates their similarity score, denoted as $E[u,v]=S_{c}(u,v)$, where $E_{i}\in\mathbb{R}^{|\mathcal{E}|}$. In this similarity graph, the similarity between two generated images is interpreted as the distance within the graph. Next, we introduce graph density [4], as a novel metric for evaluating the overall similarity of the generated batch:
Definition 3.4.
@@
If $DS(G_{i})$ is greater than the threshold $\tau$, then it is determined as a backdoor sample, otherwise, it is a clean sample and the originally generated image shall be returned to the users. The total pipeline of our method is visualized in Figure 2 and the final detection algorithm is presented in Algorithm 1.
4 Experimental Results
4.1 Experimental Settings
Attack Baselines.
@@ 4.1 Experimental Settings
To the best of our knowledge, the existing backdoor attacks on diffusion models include two unconditional-DM-based backdoor attacks: TrojDiff [11] and BadDiffusion [14], and two conditional-DM-based backdoor attacks: Rickrolling [42] and Villandiffusion [15]. We consider all four backdoor attacks as our attack baselines. It is also noted that TrojDiff supports three types of attack objectives. Detailed descriptions of the four attack methods are provided in Appendix B.
Defense Baselines.
To the best of our knowledge, there are yet methods for detecting backdoor samples on diffusion models. Moreover, existing methods are built especially for classification models and therefore not applicable to the generative models.
Models and Datasets.
@@
Different backdoor attacks are built based on different backbone models and samplers. To facilitate evaluation, TrojDiff [11] and BadDiffusion [15] are evaluated on DDPM [22], while VillanDiffusion [15] and Rickrolling [42] are evaluated on Stable Diffusion [36]. For the training datasets, we choose Cifar10 [26] and CelebA [32] for TrojDiff and BadDiffusion, and choose CelebA-HQ-Dialog [25] and Pokemon [34] for VillanDiffusion and Rickrolling.
Metrics.
@@
Following the prior works on backdoor detection, we adopt three popular metrics for evaluating the effectiveness of our detection method: Precision (P), Recall (R), and Area under the Receiver Operating Characteristic (AUC).
Implementation Details
All the models are well-trained with the default hyper-parameters reported in the original papers so that they show a good performance in generating both clean images and backdoor images. Following the previous works [27, 21], we then evaluate our detection method with a positive (i.e., attacked) and a negative (i.e., clean) dataset. For evaluations against unconditional diffusion models, we randomly generate 1000 Gaussian noises as the clean queries (negative) and construct backdoor samples (positive) accordingly by blending the trigger pattern with the Gaussian noises. For evaluations against conditional diffusion models, we split the whole dataset into 90% train and 10% test following [15]. Then, we use the textual caption in the test subset as the clean queries (negative) and construct backdoor queries (positive) accordingly. The threshold value $\tau$ is determined by a small clean held-out validation dataset, where detailed descriptions are provided in Appendix C. The pre-trained encoder is set as CLIP-ViT-B32 [35], the magnitude size is set as 4, the poisoning rate is set as 10%, and the number of validation datasets is set as 20, by default. All the hyperparameters are evaluated in the ablation studies.
4.2 Main Results
Effectiveness
@@ 4.2 Main Results
Dataset Backdoor Attacks Precision Recall AUC
Cifar10 TrojanDiff(D2I) 0.95 0.94 1.00
TrojanDiff(In-D2D) 0.93 0.93 0.98
TrojanDiff(Out-D2D) 0.93 0.92 1.00
BadDiffusion 0.93 0.95 1.00
Average 0.93 0.94 1.00
CelebaA TrojanDiff(D2I) 0.93 0.92 1.00
TrojanDiff(In-D2D) 0.90 0.89 0.96
TrojanDiff(Out-D2D) 0.91 0.92 0.98
BadDiffusion 0.97 0.95 1.00
Average 0.93 0.92 0.99
@@
Dataset Backdoor Attacks Precision Recall AUC
CelebaA-HQ-Dialog VillanDiffusion 0.92 0.95 0.96
Rickrolling 0.87 0.84 0.90
Average 0.90 0.90 0.93
Pokemon Caption VillanDiffusion 0.91 0.93 0.94
Rickrolling 0.83 0.85 0.91
Average 0.87 0.89 0.93
Table 2 presents the performance of our detection method against backdoor attacks on unconditional diffusion models, and Table 2 presents the performance on conditional diffusion models. As shown, the AUC values for different backdoor attack methods on all the evaluated datasets are over 0.9, suggesting that our method can effectively distinguish backdoor and clean samples.
Efficiency
Figure 4: Average inference speed of the proposed detection method against TrojDiff(D2I) on the Cifar10 dataset.
Figure 4 illustrates the efficiency of our detection method against TrojDiff(D2I) on the Cifar10 dataset. We query the diffusion models with 320 samples with a batch size of 64 and record the average inference speed, defined as the average time consumption for a sample. The y-axis comprises three components: ’Vanilla,’ representing the average inference speed without UFID; ’+augmented Batch Query’ representing the average inference speed after an augmented batch query; and ’+Similarity Calculation’ representing the average inference speed after similarity graph construction and calculation. Due to the increased number of query samples, UFID inevitably results in a lower inference speed than that of the vanilla mode. However, the increased time consumption is within expectations and will not significantly influence user experience in practice.
4.3 Ablation Studies
In this section, we discuss how the hyper-parameters influence the effectiveness of the detection method.
The Influence of Different Pre-trained Encoders.
Encoder $\rightarrow$ ViT-ImageNet CLIP DINO V2
ViT-B ViT-L RN50 RN50x64 ViT-B ViT-L ViT-S ViT-B ViT-L
Precision 0.94 0.95 0.89 0.85 0.91 0.80 0.81 0.85 0.80
Recall 0.93 0.95 0.88 0.81 0.91 0.79 0.70 0.78 0.65
AUC 0.98 0.99 0.88 0.95 0.97 0.88 0.99 1.00 0.99
The pre-trained encoder is a crucial component in our detection method. To evaluate its impact on the effectiveness of our detection method, we test the performance of UFID when integrated with different pre-trained encoders in Table 3. For the space limit, we only present the results against TrojDiff(In-D2D) on the Cifar10 dataset in the main manuscript. More experiments are in the Appendix D. As shown, UFID works well when integrated with different pre-trained encoders. In particular, CLIP encoders show consistently good performance across different datasets, due to their strong generalization ability from the pre-training stage.
@@ 4.3 Ablation Studies
Figure 6: Performance with different amounts of available samples.
Figure 6 investigates the impact of the magnitude size on the running-time efficiency and effectiveness, where the X-axis denotes the magnitude size, the left y-axis denotes the performance values, and the right y-axis denotes the average inference speed. The first four backdoor attacks are evaluated on the Cifar10 dataset and the remaining two backdoor attacks are evaluated on the Pokemon dataset. As the figure shows, our detection method achieves a stably satisfactory performance against all backdoor attacks when the size is over four. Additionally, a size of four would also yield a balanced trade-off on efficiency and effectiveness.
The Influence of Available Validation Dataset.
Figure 6 presents the impact of the available validation dataset on the threshold values and the performance, where the X-axis values denote the number of available validation samples and the y-axis denotes the performance values. The first four backdoor attacks are evaluated on the Cifar10 dataset and the remaining two backdoor attacks are evaluated on the Pokemon dataset. As the figure suggests, with more validation samples available, the threshold values gradually increase, and the performances tend to become more stable. However, it is also noted that if the number of validation samples becomes exceedingly large, there is a slight drop in performance. A possible explanation for this phenomenon is that more validation samples also introduce more noisy information, leading to an unexpected threshold value.
The Influence of Different Poisoning Rate.
Figure 8 explores whether the performance of UFID is sensitive to the backdoor poisoning rate. We evaluate the UFID under poisoning rate from 0.05 to 0.30. The results reveal that our method can generally provide a satisfactory performance irrelevant to the poisoning rate. Moreover, with an increase in the poisoning rate, the performance becomes more stable.
Figure 7: Performance with different poisoning rates.
Figure 8: Evaluation of UFID against adaptive attacks.
@@
We evaluate our detection method against an adaptive attacker who already has prior information about our detection method. Therefore, the attacker might try to make the generated images more diversified to avoid being detected. Specifically, rather than training a diffusion model that maps the trigger to the target images (e.g., erotic images), the attacker maps the trigger to a target domain that contains both the target images and a small number of clean images. In this way, the attacker achieves a more stealthy backdoor attack by sacrificing the attack success rate. We further define the ratio between the number of clean images to the backdoor samples in this target domain as the "blending ratio". We evaluate the UFID against TrojDiff(D2I) and employ mean square error (MSE) between the generated backdoor images and the target image (e.g., Mickey Mouse) as the attack success rate. Figure 8 presents the performance of the UFID under different blending ratios. As shown, the performance of the UFID exhibits a slight decrease when the blending ratio rises. However, the average MSE across generated backdoor samples abruptly exceeds 0.15, which suggests the failure in injecting backdoors. The right-hand side also provides some samples in the Cifar10 dataset, where we notice that images with an MSE of 0.15 to the target image are already completely different from the target image.
4.4 Discussions
Visualizations of Similarity Graphs.
@@ 4.4 Discussions
To better understand how UFID helps to detect backdoor samples, we visualize the similarity graphs in Figure 9. Due to the space limit, we only present similarity graphs against the TrojDiff In-D2D attack on the Cifar10 dataset in the main manuscript. More qualitative examples are presented in the appendix E. Each node in the similarity graph denotes the generated images of the query batch, while each edge denotes the cosine similarity scores of the embedding of any two images. As shown, the similarity scores for the clean query batch are significantly lower than those for the backdoor query batch, validating our intuitions for backdoor detections.
Visualizations of Scores Distributions.
Figure 17 and Figure 18 present the distributions of the calculated scores $\mathcal{S}_{i}$ for backdoor samples and clean samples respectively. As shown, the distribution of backdoor samples tends to be more clustered in a narrow range, while that of clean samples tends to be spread out. Moreover, there is a distinct gap between the two distributions, suggesting that our method can effectively distinguish backdoor and clean samples.
5 Related Works
Diffusion Models.
Diffusion models have become a new state-of-the-art family of generative models in image synthesis by showing strong sample quality and diversity in image synthesis [22, 38]. Since then, diffusion models have been widely applied to different types of tasks, such as image generation [39, 41], image super-resolution [37, 6], and image editing [3, 13]. Moreover, researchers also found that the representations learned by diffusion models can also be used in other discriminative tasks such as segmentation [18] and anomaly detection [33].
Backdoor Attacks and Defenses on Diffusion Models.
Recently, inspired by the great success of diffusion models, a lot of works investigate the security vulnerabilities of diffusion models by launching backdoor attacks on diffusion models. From a high-level idea, malicious backdoor attackers aim to inject a special behavior in the diffusion process such that, once the predefined trigger pattern appears on the input, the special behaviors will be activated. To achieve this goal,  [11] proposed to add an additional backdoor injection task on the training stage and maliciously alter the sampling procedure with a correction term.  [14] proposed a novel attacking strategy by only modifying the training loss function.  [1] focused on launching attacks to text-to-image tasks, by injecting backdoors into the pre-trained text encoder.  [15] proposed a unified framework that covers all the popular schemes of diffusion models, including conditional diffusion models and unconditional diffusion models. However, backdoor defenses on diffusion models are highly under-explored. To the best of our knowledge, only  [1] investigated backdoor defenses on diffusion models. However, their work aims to detect whether a given model is backdoored, while our tasks focus on filtering backdoor samples for diffusion models in the inference stage, which are fundamentally different.
6 Conclusion and Future Directions
@@
In this paper, we propose a simple unified framework for backdoor detection on diffusion models under the MaaS setting. Our framework is first motivated by a strict causality analysis on image generation and further validated by theoretical analysis. Motivated by the analysis, we design a unified method for distinguishing backdoor and clean samples for both conditional and unconditional diffusion models. Extensive experiments demonstrate the effectiveness of our method. Despite the great success, there are still many directions to be explored in the future. Firstly, our method relies heavily on a pre-trained encoder, which is not always practical. It is promising to design some model-free metrics to detect backdoor samples on diffusion models. Secondly, extending our unified framework to the newly emerging diffusion models is also of great interest.
Impact Statment

### v2

@@ Introduction
Conditional diffusion models accept input from other modalities to guide the generation of user-intended images. For instance, in stable diffusion (Rombach et al. 2021), textual input is used to query the model. However, it is evident that the detection approach employed for unconditional diffusion models cannot be directly applied to conditional diffusion models due to the inability to introduce Gaussian noise to discrete textual input. To adapt our detection method for conditional diffusion models, we have extended the detection approach with slight modifications, as depicted in the second row of Figure 2. We leverage variations in output diversity to distinguish between clean and backdoor samples. To further enlarge this diversity gap and enhance distinguishability in conditional setting, we propose appending the input text $x_{i}$ with a random phrase $ph_{j}$ selected from a diverse phrase pool containing completely distinct phrases, such as ”Iron Man” and ”Kitchen Dish Washer,” denoted as =$\mathbb{N}=\{ph_{1},ph_{2},...ph_{|\mathbb{N}|}\}$. For each input $x_{i}$, we repeat this process $|\mathbb{M}|$ times, where $|\mathbb{M}|\ll|\mathbb{N}|$. This results in an augmented input batch $\mathbb{I}=\{x_{i}\}\cup\{x_{i}^{j}|x_{i}^{j}=x_{i}\oplus ph_{j},\forall 1\leq j\leq|\mathbb{M}|\}$, where $\oplus$ denotes the string appending operator. Similarly, we can generate an image batch $y_{i}$ by querying the target stable diffusion model using Equation 3.
A Unified Framework for Backdoor Detection
@@
As previously discussed, the diversity of images generated by a specific input $i$ can be used to detect backdoors. To quantify this diversity, we employ a two-step process that involves calculating both pairwise and overall similarities within the generated batch. The detailed steps are as follows:
@@
Pairwise Similarity Calculation. Initially, we calculate semantic embeddings of generated images through a pre-trained image encoder (e.g., ViT-ImageNet (Wu et al. 2020) and CLIP (Radford et al. 2021)), denoted as $f_{\mathcal{E}}(\cdot)$. Subsequently, we calculate the local similarity for each pair of images in the generated batch using cosine similarity, represented by $S_{c}(\cdot,\cdot)$.
@@
Graph Density Calculation. Following this, we construct a weighted graph and compute its graph density to represent the overall similarity of all images within the graph. Specifically, let $G_{i}=(V_{i},\mathcal{E}_{i})$ represent the similarity graph for input sample $x_{i}$, where $|V_{i}|=|\mathbb{M}|$ constitutes the set of vertices (symbolizing the generated images) and $\mathcal{E}_{i}$ is the set of edges. Each edge’s weight, connecting a pair of images $u,v\in V$, indicates their similarity score, denoted as $E[u,v]=S_{c}(u,v)$, where $E_{i}\in\mathbb{R}^{|\mathcal{E}|}$. In this similarity graph, the similarity between two generated images is interpreted as the distance within the graph. Next, we introduce graph density (Balakrishnan and Ranganathan 2012), as a novel metric for evaluating the overall similarity of the generated batch:
Definition 4.
@@
If $DS(G_{i})$ is greater than the threshold $\tau$, then it is determined as a backdoor sample, otherwise, it is a clean sample and the originally generated image shall be returned to the users. The total pipeline of our method is visualized in Figure 2 and the final detection algorithm is in Algorithm 1.
Remark 5 (Reliance on Pre-trained Encoders).
Relying on a pre-trained encoder $f_{\mathcal{E}}(\cdot)$ might not always be feasible in practice. To relax this assumption, we also explore using a model-free metric structural similarity index measure (SSIM) to calculate the pairwise similarity between images. We report the evaluation results in the Appendix.
Remark 6 (Applicability of UFID).
The effectiveness of UFID is based on a practical assumption that backdoor generations are more similar than clean generations, which has been implicitly made in the previous backdoor attacks work, e.g., the backdoor generations share a similar style (Struppek, Hintersdorf, and Kersting 2022; Huang, Guo, and Juefei-Xu 2023), object (Chou, Chen, and Ho 2023b, a), or semantic concept (Chen, Song, and Li 2023). It is also observed that when the target images of backdoor attacks are as diversified as the clean images (Zhai et al. 2023), the performance of UFID will be limited. However, we could incorporate supplementary information to enhance the UFID. Detailed evaluations are provided in the Appendix.
Experiments
Experimental Settings
@@
Attack Baselines. To the best of our knowledge, the existing backdoor attacks on diffusion models include two unconditional-DM-based backdoor attacks: TrojDiff (Chen, Song, and Li 2023) and BadDiffusion (Chou, Chen, and Ho 2023a), and three conditional-DM-based backdoor attacks: Rickrolling (Rick) (Struppek, Hintersdorf, and Kersting 2022), Villandiffusion (VillanDiff) (Chou, Chen, and Ho 2023b), and Personalization (Personal) (Huang, Guo, and Juefei-Xu 2023). We consider all five backdoor attacks as our attack baselines. It is also noted that TrojDiff, Rickrolling, and Personalization all support diversity-preserving backdoor attacks, where the attackers’ target images are diversified. Detailed descriptions of them are provided in the Appendix.
Defense Baselines. We compared our method with the existing well-established defense method TERD (Mo et al. 2024) on the unconditional diffusion models and Shield (Wang et al. 2024b) on the conditional diffusion models. It is noted that both TERD and Shield require additional white-box access to the model weights and structures, which are not always feasible in our MaaS setting.
@@
Models and Datasets. Different backdoor attacks are built based on different backbone models and samplers. To facilitate evaluation, TrojDiff and BadDiffusion are evaluated on DDPM, while VillanDiffusion, Rickrolling, and Personalization are evaluated on Stable Diffusion v1.4 (Rombach et al. 2021). For the training datasets, we choose CIFAR10 (Krizhevsky, Nair, and Hinton ) and CelebA (Liu et al. 2015) for TrojDiff and BadDiffusion, and choose CelebA-D (Jiang et al. 2021) and Pokemon (Pinkney 2022) for VillanDiffusion, Rickrolling, and Personalization.
@@
Metrics. Following the prior works on backdoor detection, we adopt three popular metrics for evaluating the effectiveness of our detection method: Precision (P), Recall (R), and Area under the Receiver Operating Characteristic (AUC).
Implementation Details. All the models are well-trained with the default hyper-parameters reported in the original papers. Following the previous works (Lee et al. 2018; Guo et al. 2023), we evaluate our detection method with a positive (i.e., attacked) and a negative (i.e., clean) dataset. Due to space limits, the details for constructing the two datasets and the default hyper-parameters are provided in the Appendix.
Main Results
Effectiveness.
UFID(black-box) TERD(white-box)
Dataset Attacks P R AUC P R AUC
Cifar10 TrojDiff(D2I) 0.95 0.94 1.00 1.00 1.00 1.00
TrojDiff(In) 0.93 0.93 0.98 1.00 1.00 1.00
TrojDiff(Out) 0.93 0.92 1.00 1.00 1.00 1.00
BadDiffusion 0.93 0.95 1.00 1.00 1.00 1.00
Average 0.93 0.94 1.00 1.00 1.00 1.00
CelebaA TrojDiff(D2I) 0.93 0.92 1.00 1.00 1.00 1.00
TrojDiff(In) 0.90 0.89 0.96 1.00 1.00 1.00
TrojDiff(Out) 0.91 0.92 0.98 1.00 1.00 1.00
BadDiffusion 0.97 0.95 1.00 1.00 1.00 1.00
Average 0.93 0.92 0.99 1.00 1.00 1.00
@@
UFID(black-box) Shield(white-box)1
Dataset Attacks P R AUC P R AUC
CelebaA-D VillanDiff 0.92 0.95 0.96 0.80 0.95 -
Rick(TPA) 0.87 0.84 0.90 0.96 0.85 -
Rick(TAA) 0.82 0.81 0.87 - - -
Personal 0.76 0.72 0.82 - - -
Average 0.85 0.84 0.89 0.88 0.90
Pokemon VillanDiff 0.91 0.93 0.94 - - -
Rick(TPA) 0.83 0.85 0.91 - - -
Rick(TAA) 0.80 0.81 0.87 - - -
Personal 0.73 0.77 0.81 - - -
Average 0.82 0.85 0.89 - - -
  • 1
    Due to unavailable codes, we use the reported values in the original paper directly.
@@
Table 1 presents the performance of our detection method against backdoor attacks on unconditional diffusion models, while Table 2 presents the performance on conditional diffusion models. As shown, the AUC values for different backdoor attack methods on all the evaluated datasets are over 0.8, suggesting that our method can effectively distinguish backdoor and clean samples. Compared to the baseline methods, UFID shows a comparable performance with a slight drop. However, the decrease is reasonable as the two baselines are white-box methods, which require additional access to the model weights and structures. The similar performance drop has been observed in the previous studies (Guo, Li, and Liu 2021; Guo et al. 2023).
Figure 3: Average inference speed against TrojDiff(D2I) on the Cifar10.
Figure 4: Average inference speed against VillanDiffusion on the Pokemon dataset.
Efficiency. Figure 3 illustrates the efficiency of our detection method against TrojDiff(D2I) on CIFAR10 (32$\times$32$\times$ 3) dataset. We query the diffusion models with 320 samples with a batch size of 64 and record the average inference speed, defined as the average time consumption for a sample. The y-axis comprises three components: ’Vanilla,’ representing the average inference speed without UFID; ’+augmented Batch Query’ representing the average inference speed after an augmented batch query; and ’+Similarity Calculation’ representing the average inference speed after similarity graph construction and calculation. Due to the increased number of query samples, UFID inevitably results in a lower inference speed than that of the vanilla mode. However, the increased time consumption is within expectations. Figure 4 presents the efficiency of the UFID against VillanDiffusion on the Pokemon (512$\times$512$\times$3) dataset. Due to the acceleration techniques in the modern sampler (Song, Meng, and Ermon 2020; Lu et al. 2022), stable diffusion models are usually denoised with only a small number of steps (e.g., 25) to generate high-resolution images. Therefore, we could observe that the UFID only slightly increases the inference time. To further investigate how the selected inference steps influence the efficiency, we evaluate the inference speed with different numbers of inference steps. The results show that when setting the inference step from 10 to 75, the inference overhead is acceptable.
Ablation Studies
In this section, we discuss how the hyper-parameters influence the effectiveness of the UFID.
The Influence of Different Pre-trained Encoders. The pre-trained encoder is important in our detection method. To evaluate its impact on the effectiveness of our detection method, we test the performance of UFID when integrated with different pre-trained encoders. For the space limit, we present the results in the Table 3 - Table 8 in the Appendix. According to the tables, UFID works well when integrated with different pre-trained encoders. In particular, CLIP encoders show consistently good performance across different datasets, due to their strong generalization ability from the pre-training stage.
@@
Figure 5 investigates the impact of the magnitude size on the running-time efficiency and effectiveness, where the left y-axis denotes the performance values, and the right y-axis denotes the average inference speed. Due to the space limit, we only present partial results in the main manuscript, while the remaining parts can be found in the Appendix. As the figure shows, our detection method achieves a stably satisfactory performance against all backdoor attacks when the size is over four. Additionally, a size of four would also yield a balanced trade-off on efficiency and effectiveness.
The Influence of Available Validation Dataset. Figure 16 and Figure 17 in the Appendix presents the impact of the available validation dataset on the performance, where the X-axis values denote the number of available validation samples and the y-axis denotes the performance values. As the figure suggests, with more validation samples available, the performances tend to become more stable. However, it is also noted that if the number of validation samples becomes exceedingly large, there is a slight drop in performance. A possible explanation for this phenomenon is that more validation samples also introduce more noisy information, leading to an unexpected threshold value.
The Influence of Image Size. Figure 20 in the Appendix investigates whether UFID’s performance will be influenced when handling high-resolution images. Specifically, we evaluate UFID against TrojDiff and BadDiffusion on an augmented CIFAR10 dataset, where the image size is manually scaled to 64, 128, and 256. We could see that the image size does not have any influence on the performance, demonstrating UFID’s potential to handle high-resolution images.
The Influence of Different Poisoning Rate. Figure 15 in the Appendix explores whether the performance of UFID is sensitive to the backdoor poisoning rate. We evaluate the UFID under poisoning rate from 0.05 to 0.30. The results reveal that our method can perform satisfactorily under different poisoning rates. Moreover, with the increase of the poisoning rate, the performance becomes more stable.
Discussions
Visualizations of Similarity Graphs. To better understand how UFID helps to detect backdoor samples, we visualize the similarity graphs in Figure 6. Due to the space limit, we only present similarity graphs against the TrojDiff(In-D2D) on CIFAR10 in the main manuscript. More qualitative examples are presented in the appendix. Each node in the similarity graph denotes the generated images of the query batch, while each edge denotes the cosine similarity scores of the embedding of any two images. As shown, the similarity scores for the clean query batch are significantly lower than those for the backdoor query batch, validating our intuitions for backdoor detection.
Visualizations of Scores Distributions. Figure 18 and Figure 19 in the Appendix present the distributions of the graph density $\mathcal{S}_{i}$ for backdoor samples and clean samples respectively. As shown, the distribution of backdoor samples tends to be more clustered in a narrow range, while that of clean samples tends to be spread out. Moreover, there is a distinct gap between the two distributions, suggesting that UFID can effectively distinguish backdoor and clean samples.
@@
We evaluate our detection method against an adaptive attacker who already has prior information about our detection method. Therefore, the attacker might try to make the generated images more diversified to avoid being detected. Specifically, rather than training a diffusion model that maps the trigger to the target images (e.g., erotic images), the attacker maps the trigger to a target domain that contains both the target images and a small number of clean images. In this way, the attacker achieves a more stealthy backdoor attack by sacrificing the attack success rate. We further define the ratio between the number of clean images to the backdoor samples in this target domain as the ”blending ratio”. We evaluate the UFID against TrojDiff(D2I) and employ mean square error (MSE) between the generated backdoor images and the target image (e.g., Mickey Mouse) as the attack success rate. Figure 21 in the Appendix presents the performance of the UFID under different blending ratios. As shown, the UFID’s performance gradually decreases when the blending ratio rises. However, the average MSE across generated backdoor samples abruptly exceeds 0.15, which suggests the failure in injecting backdoors. The right-hand side also provides some samples in the CIFAR10 dataset, where we notice that images with an MSE of 0.15 to the target image are already completely different from the target image.
Conclusion and Future Directions
@@
In this paper, we propose a simple unified framework for backdoor detection on diffusion models under the MaaS setting. Our framework is first motivated by a causality analysis on image generation and further validated by theoretical analysis. Motivated by the analysis, we design a unified method for distinguishing backdoor and clean samples for both conditional and unconditional diffusion models. Extensive experiments demonstrate the effectiveness of our method. Despite the great success, there are still many directions to be explored in the future. For example, UFID still requires a small amount of validation samples to determine the threshold. Can we relax the assumption?
Acknowledgments
The work is supported in part by the U.S. Office of Naval Research Award under Grant Number N00014-24-1-2668 and the National Science Foundation under Grants IIS-2316306, CNS-2330215, CCF-1918656, CNS-2317193, IIS-2331315.